OpenStudy (anonymous):

How do I find the radius of v = pi*r^2*h

OpenStudy (tyteen4a03):

This is the equation of a circular cylinder, where r is the radius of the base circle.

OpenStudy (anonymous):

\[v = \pi{r^2}h\] v = 5m^3 and h = 1.5m

OpenStudy (tyteen4a03):

@joshhhunt Just plug in the values and solve. :)

OpenStudy (anonymous):

\[\frac{ V }{ \pi h } = r^{2}\] Take the square root of both sides and you'll get the radius.

OpenStudy (anonymous):

How do I go about getting the square root of both sides?

OpenStudy (anonymous):

Ok so I have 125 = \[125 = \pi \times r^2 \times 1.5m\] I still don't understand how I find out the value of r

OpenStudy (anonymous):

Look at my formula. You have the volume V and the height h. Input them: 125 / (Pi * 1.5) then take the square root of that: \[\sqrt{\frac{ 125 }{ 1.5 \pi }}\]

OpenStudy (anonymous):

So I have 5.568... Am I right or have I done it wrong?

OpenStudy (anonymous):

Wait no I cubed it. Actually it should be 3.141...?

OpenStudy (anonymous):

sqrt(125 / (1.5 * pi)) ~ 5.15 Remember to use the parentheses. ; )